OpenStudy (anonymous):

which transformation of the graph of y=x^2 would result in the graph of y=x^2+2

OpenStudy (anonymous):

Is this a multiple choice question?

OpenStudy (anonymous):

yeah here are the answers: a) r(y-2) b) T(0,2) c) R(0,90degrees) d) Dilation 2

OpenStudy (jdoe0001):

\(\large { \textit{function transformations} \\ \quad \\ \begin{array}{llll} \begin{array}{llll} shrink\ or\\ expand\\ by\ {\color{purple}{ A}}\cdot {\color{blue}{ B}}\end{array} \qquad \begin{array}{llll} vertical\\ shift\\ by \ {\color{green}{ D}} \end{array} \begin{array}{llll}{\color{green}{ D}} > 0& Up\\ {\color{green}{ D}} < 0 & Down\end{array} \\ \qquad \downarrow\qquad\qquad\qquad \downarrow\\ y = {\color{purple}{ A}} ( {\color{blue}{ B}}x + {\color{red}{ C}} ) + {\color{green}{ D}}\\ \qquad\qquad\uparrow \\ \qquad\begin{array}{llll} horizontal\\ shift\\ by \ \frac{{\color{red}{ C}}}{{\color{blue}{ B}}}\end{array} \begin{array}{llll}\frac{{\color{red}{ C}}}{{\color{blue}{ B}}} > 0 & to \ left\\ \frac{{\color{red}{ C}}}{{\color{blue}{ B}}} < 0& to \ right\end{array} \end{array} }\) what do yiou think?

OpenStudy (anonymous):

A rotation would not move the line 2 spaces up. A dilation would extend the slope of the line, not move it up. A reflection over y=2 would cause the slope to be negative. So, what answer are you left with?

OpenStudy (anonymous):

oops dont wanna ruin his choice :D

OpenStudy (anonymous):

that makes sense, thank u @BunnyBree !!

OpenStudy (anonymous):

not really what this means @jdoe0001 but thanks anyway!

OpenStudy (anonymous):

You're very welcome! (: